p=8,000 at 6.5% compounded daily for 15 years. find future value of given principal p, and find the interest earned in the given period
Answer by stanbon(75887)

p=8,000 at 6.5% compounded daily for 15 years. find future value of given principal p, and find the interest earned in the given period
----
A(15) = 8000(1+(0.065/365))^(365*15)
-----
A(15) = 8000(1.0001781)^(5475)
----
A(15) = 8000*2.6512
----
A(15) = $21209.56
======================
Cheers,
Stan H.
========================